OpenStudy (anonymous):

Solve for x. 3x2 – 75 = 0 A. {25, –25} B. {5, –5} C. {5} D. {25}

OpenStudy (anonymous):

3x^2=75 x^2=75/3 x^2=25 x=sqreroot 25 which is +5,-5 so ans is B

OpenStudy (anonymous):

No answer of your question is correct.. and @madrockz have given correct explaination

OpenStudy (anonymous):

THANKS @bhaskar.pathk

OpenStudy (anonymous):

@TakeA.I.M. : keeping -5 in the given equation.. 3x2 – 75 = 0 3(-5)^2 -75 3 X 25 -75= 0 as square of 5 and -5 are equal ie 25 hope you get it..